OpenStudy (ellaella02):

Which equation shows the substitution method being used to solve the system of linear equations? x + y = 6 x = y + 5 A. x = (x – 6) + 5 B. (y + 5) + y = 6 C. x + (y + 5) = 6 D. x + y = y + 5

OpenStudy (alexandervonhumboldt2):

substitute the value of x into the first equation: x + y = 6 x = y + 5 => (y+5)+y=6 => 2y=1 => y=1/2 x+1/2=6 x=5.5 thus y=0.5 and x=5.5

OpenStudy (ellaella02):

Thank you so much (:

OpenStudy (ellaella02):

But none of them are right?... @AlexandervonHumboldt2

OpenStudy (alexandervonhumboldt2):

no you need to susbtitute the value of x into first equation. x=y+5. substitue this instead of x: (y+5)+y=6